# Local algebraicity and localization of the Bergman kernel on Stein spaces with finite type boundaries

### v1

@@ 3. The type-degree estimate: Proof of Theorem 1
We may take any $c\geq n+s+1$. In the two dimensional case, $\lambda$ can be taken as $2^{-m}$ (by [Koh72]) and $\leq 1/m$ in higher dimensions (by [Cat87b, Theorem 3.3]), where $m$ is the type of the boundary point $p$.
@@
By the work of Catlin ( [Cat83], [Cat87b]), having a finite type boundary (in the sense of D’Angelo) is the necessary and sufficient condition for the following subelliptic estimate. Although in the original works of Catlin, this estimate is established for smoothly bounded, pseudoconvex domains with finite type boundary in $\mathbb{C}^{n}$, we note that the estimate is purely local in nature and, hence, the same holds for such domains in complex manifolds as well. For a smooth form $u\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$ which is compactly supported in $U\cap\overline{M}$, we have:
@@
We now proceed to prove Claim A.1. We observe that if $g$ is a smooth form in $\overline{M}$ then so is $Ng$. This follows from the Sobolev regularity gain property of the operator $N$; see Proposition C.1; see also [Koh72, Theorem 6.3], [Cat87b], [Str10, Theorem 3.4], [DK99, Theorem 3.5]. By definition of the $\bar{\partial}$-Neumann operator $N$, we have $Ng\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$, ¯$\bar{\partial}Ng\in Dom(\bar{\partial}^{*})$ and ¯$\bar{\partial}^{*}Ng\in Dom(\bar{\partial})$. Since $\xi\in C^{\infty}(\overline{M})$, we further have $\xi Ng\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$ (cf. [Str10, page 11-12] and [CS01, page 106]). Thus, we may apply (A.8) to $u=\xi Ng$, which gives us, for all $s\in\mathbb{N}\cup\{0\}$,
@@ Appendix C Proof of continuity of $N$ on the $H_{s}^{\dagger}(M)$ Sobolev spaces
The second inequality follows because the $\bar{\partial}$-Neumann operator $N$ is a bounded operator from $L^{2}(M)$ to itself. The third inequality follows as $\Box N=id-H$, where $id$ is the identity operator and $H$ is the $L^{2}$ projection operator to the $L^{2}$ harmonic forms. The sixth inequality above follows as $Hg$ is a harmonic form and, as proved in Claim C.1, the space of harmonic forms is finite dimensional and, hence, all norms on it are equivalent. The seventh inequality follows as $H$ is a $L^{2}$ projection operator. According to the definition of the Sobolev norms on $M$ in Section 2, the above proves the first part of Proposition C.1. For proving the second part, we recall that, for $s>0$, we have defined $|\!|{}u|\!|{}_{-s}^{\dagger}=\sup_{g\in C^{\infty}_{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})}|(u,g)|/|\!|{}g|\!|{}_{s}$. For any $g\in C^{\infty}_{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*}),$ we note that, for $s>0$,

### v2

@@ 3. The type-degree estimate: Proof of Theorem 1
Remark 3.1.
In the proof above we applied our localization result, Theorem 3 to obtain a Hsiao-Savale type expansion for the Bergman kernel of the Stein domain $\Omega$, see (3.4). It is possible that the asymptotic expansion of Hsiao-Savale in [HS22] applies directly in this Stein case with an adjustment of their proof. The authors have not tried to verify this.
@@
We may take any $c\geq n+s+1$. In the two dimensional case, $\lambda$ can be taken as $2^{-m}$ (by [Koh72]) and $\leq 1/m$ in higher dimensions (by [Cat87, Theorem 3.3]), where $m$ is the type of the boundary point $p$.
@@
By the work of Catlin ( [Cat83], [Cat87]), having a finite type boundary (in the sense of D’Angelo) is the necessary and sufficient condition for the following subelliptic estimate. Although in the original works of Catlin, this estimate is established for smoothly bounded, pseudoconvex domains with finite type boundary in $\mathbb{C}^{n}$, we note that the estimate is purely local in nature and, hence, the same holds for such domains in complex manifolds as well. For a smooth form $u\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$ which is compactly supported in $U\cap\overline{M}$, we have:
@@
We now proceed to prove Claim A.1. We observe that if $g$ is a smooth form in $\overline{M}$ then so is $Ng$. This follows from the Sobolev regularity gain property of the operator $N$; see Proposition C.1; see also [Koh72, Theorem 6.3], [Cat87], [Str10, Theorem 3.4], [DK99, Theorem 3.5]. By definition of the $\bar{\partial}$-Neumann operator $N$, we have $Ng\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$, ¯$\bar{\partial}Ng\in Dom(\bar{\partial}^{*})$ and ¯$\bar{\partial}^{*}Ng\in Dom(\bar{\partial})$. Since $\xi\in C^{\infty}(\overline{M})$, we further have $\xi Ng\in Dom(\bar{\partial})\cap Dom(\bar{\partial}^{*})$ (cf. [Str10, page 11-12] and [CS01, page 106]). Thus, we may apply (A.8) to $u=\xi Ng$, which gives us, for all $s\in\mathbb{N}\cup\{0\}$,
@@ Appendix C Proof of continuity of $N$ on the $H_{s}^{\dagger}(M)$ Sobolev spaces
The second inequality follows because the $\bar{\partial}$-Neumann operator $N$ is a bounded operator from $L^{2}(M)$ to itself. The third inequality follows as $\Box N=id-H$, where $id$ is the identity operator and $H$ is the $L^{2}$ projection operator to the $L^{2}$ harmonic forms. The sixth inequality above follows as $Hg$ is a harmonic form and, as proved in Claim C.1, the space of harmonic forms is finite dimensional and, hence, all norms on it are equivalent. The seventh inequality follows as $H$ is an $L^{2}$ projection operator. According to the definition of the Sobolev norms on $M$ in Section 2, the above proves the first part of Proposition C.1. For proving the second part, we recall that, for $s>0$, we have defined $|\!|{}u|\!|{}_{-s}^{\dagger}=\sup_{g\in C^{\infty}_{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*})}|(u,g)|/|\!|{}g|\!|{}_{s}$. For any $g\in C^{\infty}_{p,q}(\overline{M})\cap Dom(\bar{\partial}^{*}),$ we note that, for $s>0$,